OpenStudy (anonymous):

use spherical coordinates to evaluate the integral S[-1,1]S[sqrt(1-x^2), 0] S[sqrt(1-x^2-y^2),0] e^-(x^2+y^2+z^2)^(3/2)

OpenStudy (turingtest):

do you know the formula to conversion from cartesian coordinates to spherical ones?

OpenStudy (anonymous):

i know rho squared = x^2 + y^2 + z^2 but i dont know how to convert the bounds

OpenStudy (turingtest):

identify the shape you are integrating over, what is it?

OpenStudy (anonymous):

would theta be 0 to pi and rho be from 0 to 1?

OpenStudy (anonymous):

its a cone?

OpenStudy (turingtest):

not is is a sphere not sure what you mean by theta, some people use theta and phi intechangeably byt your bounds for rho are correct

OpenStudy (turingtest):

no, it is...

OpenStudy (turingtest):

*

OpenStudy (turingtest):

\[0\le z\le\sqrt{1-x^2-y^2}\implies z^2+y^2+x^2=1\]which is a sphere of radius 1

OpenStudy (anonymous):

yeah I'm stuck as to how i would change the bounds since dv = rho^2 sin(phi) d(rho) d(phi) d(theta)

OpenStudy (turingtest):

dV is not really waht we look at in changing the bounds sorry, brb in the meantime think about what the bounds would be to represent a sphere

OpenStudy (anonymous):

x=1 y=1 z = 1 all of them are plus and minus

OpenStudy (turingtest):

I mean in terms of rho, theta, and phi

OpenStudy (anonymous):

theta would be 0 to pi and phi would be 0 to pi/4?

OpenStudy (anonymous):

wait would theta be 0 to 2pi?

OpenStudy (turingtest):

theta is zero to 2pi and rho is from 0 to pi I think let me double-check when I get off the phone

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

whoops the bounds are messed up the corrected ? is use spherical coordinates to evaluate the integral S[-1,1]S[0, sqrt(1-x^2)] S[0,sqrt(1-x^2-y^2)] e^-(x^2+y^2+z^2)^(3/2)

OpenStudy (turingtest):

oh I figured that much

OpenStudy (anonymous):

so rho is 0 to 1 and theta is from 0 to 2pi right? and phi's what we're missing

OpenStudy (turingtest):

if we are doing the whole sphere phi must be either -pi/2 to pi/2 or 0 to pi (I actually don't think it makes a difference, but I may have forgotten that detail)

OpenStudy (anonymous):

i dont think it makes a difference since the range ends up being the same

OpenStudy (turingtest):

I agree

OpenStudy (anonymous):

so changing the values would be S[0,2pi]S[0,pi]S[0,1] rho^2 sin(phi) d(rho) d(phi) d(theta)?

OpenStudy (anonymous):

well S[0,2pi]S[0,pi]S[0,1] (e^-(rho^3))* rho^2 sin(phi) d(rho) d(phi) d(theta)

OpenStudy (turingtest):

right, which is pretty easy to integrate, eh?

OpenStudy (anonymous):

lol right...

OpenStudy (turingtest):

I think we got it then. Good luck!

OpenStudy (anonymous):

thanks

OpenStudy (turingtest):

welcome

OpenStudy (anonymous):

hey on my ? isn't it a quarter circle due to the sqrts so shouldn't the limits be 0 ≤ Θ ≤ π, 0 ≤ φ ≤ π/2, and 0 ≤ ρ ≤ 1? thanks

OpenStudy (turingtest):

if we take only the positive sqrt it would be just the upper half of the sphere\[0\le\theta\le2\pi\]and\[0\le\phi\le\frac\pi2\]rereading I think you are actually right, and that we are only doing the upper half

OpenStudy (turingtest):

I think example2 here is similar: http://tutorial.math.lamar.edu/Classes/CalcIII/TISphericalCoords.aspx